The ratio of male students to female students at a certain university is 4 to 5. If there is a total of 9,450 students, find the number of male students and the number of female students.

Answer by stanbon(75887)

The ratio of male students to female students at a certain university is 4 to 5. If there is a total of 9,450 students, find the number of male students and the number of female students.
# of males:: 4x
# of females:: 5x
---------------
4x + 5x = 9450
9x = 9450
x = 1050
------
# of males = 4x = 4*1050 = 4200
# of females = 5x = 5*1050 = 5250
---------------
Cheers,
Stan H.
------------